Solve the equation for x and show all your work.

6+ (x –5)/2=(x –7)/3

Please help me solve this problems and each step in order to understand it.
Thank you in advance.

Answer by Cromlix(4381)

First of all clear the brackets
6 + (x - 5)/2 => 6 + x -5/2 => x + 1/2
So the equation becomes:-
x + 1/2 = x - 7/3
To remove the lower numbers(2 and 3)
you find one number that both 2 and 3
will divide into. This is 6
2 divides into 6 then 3 goes up to
multiply x + 1 = 3x + 3
3 divides into 6 then 2 goes up to
multiply x - 7 = 2x - 14.
Now as both equations have the same denominator (6)
so we can equal one to the other.
3x + 3 = 2x - 14
Sort out like terms
3x - 2x = - 14 - 3
x = - 17
To check put x = -17 into:-
x + 1/2 => -17 + 1/2 = -8
x - 7/3 => -17 - 7/3 = -8 Both agree